Circular No. 7735

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ex IN UGC 3595 G. M. Hurst, Basingstoke, England, reports the discovery of an ap-

parent supernova (mag 17.2) by T. Boles, Coddenham, England, on a 60-s unfiltered CCD frame (limiting mag 18.5) taken with a 0.36-m telescope on Oct. 16.195 UT in the course of the U.K. Nova/Supernova Patrol. Boles obtained a confirming image on Oct. 17.096, which shows SN 2001ex at mag 17.1. The new object is located at α = 6^{h}55^{m}59^{s}.89, δ = +55^{o}24′03′′.0 (equinox 2000.0; average from measures of the two images), which is 5′′ west of the nucleus of its host galaxy. A frame taken by Boles on Sept. 16.075 shows nothing at this position (limiting mag 18.5). Palomar Schmidt blue and red images from 1993 and 1989 also show nothing at this location.

SUPERNOVA 2001ey IN MCG –01-57-10 T. Puckett and M. Marcus, Mountain Town, GA, report the discovery

of an apparent supernova (mag 17.6) on an unfiltered CCD frame (limiting mag 19.0) taken with the Puckett Observatory 0.35-m automated patrol telescope on Oct. 16.17 UT. SN 2001ey is located at α = 22^{h}26^{m}30^{s}.77, δ = −6^{o}23′28′′.6 (equinox 2000.0), which is 5′′.7 east and 10′′.4 north of the center of MCG –01-57-10. The new object was confirmed on CCD frames taken on Oct. 17.14. SN 2001ey is not present on images taken by Puckett on 1999 Aug. 8 and 18 (limiting mag ∼ 20.0) or on Palomar Sky Survey images taken on 1987 Sept. 15 (limiting mag ∼ 21.0) and 1983 July 16 (limiting mag ∼ 19.7).

2001 OE_{84}

P. Pravec and P. Kušnirák, Ondřejov Observatory, write: “Our photometric observations of this Amor-type minor planet (cf. MPEC 2001-P09, 2001-T04) on Oct. 15.1 and 16.0 UT reveal a rotation period of 29.190 ± 0.002 min, indicating that it is a body with non-zero tensile strength (monolithic), the first known minor planet of this kind in the km-size range (H = 17.7, diameter ∼ 0.9 km). Its lightcurve and further information are

posted at http://sunkl.asu.cas.cz/~ppravec/2001oe84.htm. Observa-

tions with other techniques (e.g., spectral, infrared) are desirable.”

2001 October 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green